OpenStudy (anonymous):

Find a polynomial with integer coefficients, with leading coefficient 1, degree 5, zeros i and 7- i, and passing through the origin P(x)=

Directrix (directrix):

zeros i and 7- i, Non-real roots appear in conjugate pairs. If i is a root, then so is -i If 7- i is a root, then so is 7 + 1 That is a total of 4 roots. If the polynomial is of degree 5, there is a fifth root. Because the polynomial passes through the origin, then x = 0 is the fifth root.

Directrix (directrix):

If x = 0 is a root, then x - 0 or x is a factor. If x = i is a root, then x - i is a factor. If x = -i is a root, then x + i is a factor. If x = 7 - i is a root, then x - 7 + i is a factor. If x = 7 + i is a root, then x - 7 - i is a factor.

Directrix (directrix):

The polynomial is the product of these 5 factors. P(x) = x * (x - i ) * (x + i) * (x - 7 + i) * ( x - 7 - i) Your Task: Multiply those factors and simplify. @aprilrod23 Then, there's the matter of integer coefficient.

Directrix (directrix):

@aprilrod23 Let an online program do all that multiplication if that is allowed. Go to wolframalpha.com and enter: x(x-i) (x+i)(x -7+i)( x-7-i)= After the "Wolf" calculates, scroll down and look at the alternate forms of writing the product. Post what you get, okay?

Directrix (directrix):

@aprilrod23 --> is just looking around

OpenStudy (anonymous):

thank you I got the answer